There are 3 bonds in this portfolio

  1. A 20 year semi-annual bond where:
    1. Par is $100,000
    2. C = 4%
    3. r = 6%         
  2. A 30 year annual bond where:
    1. Par is $50,000
    2. C = 3%
    3. r = 3.5%
  3. a 5 year semi-annual bond where:
    1. Par is $30,000
    2. C = 3%
    3. r = 2.25%

Find the value of each bond in the portfolio & how much is the portfolio worth?
Answer #1

years semi annual bond a par value (A) years period (N) coupon coupon rate rate 100,000.00 20 40 4% 2000 6.00% 3.00% half yea

30,000.00 years semi annual Bond C par value (A) years period (N) coupon coupon rate rate 10 3% 450 2.50% 1.25% hafl yearly y

FORMULAE

years semi annual bond a par value (A) years period (N) coupon coupon rate rate 100000 20 =+B13*2 0.04 =+B12*B15/2 0.06 =+B17

Bond C par value (A) 30000 semi annual period (N) coupon coupon rate rate =+B34*2 0.03 =+B33*B36/2 0.025 =+B38/2 hafl yearly

please ask if any query.